# Development of a comprehensive PMT optical model for the JUNO experiment

### v1

@@ 3.1 PDE data
For each 20-inch PMT, the PDE is measured using a container system that employs a 420 nm light source to illuminate the entire PMT PC in air. Additionally, a subset of PMTs is evaluated for the PDE uniformity test using a scanning station system, which scans the PC with seven LEDs that provide 420 nm light nearly perpendicular to the surface at seven fixed zenith angles, meanwhile the LEDs can be rotated to cover the different azimuth angles. The PDE obtained from the container system is then corrected to align with the PDE as defined by the scanning station system, which averages the PDE values obtained at the positions of the seven LEDs, weighted by their respective surface areas, and Table 1 summarizes the surface area weights along with the zenith angles for each LED for both NNVT and HPK PMTs. In this study, the PDE values averaged over azimuthal angles for seven zenith angles are utilized to extract the thickness maps of the PC and ARC.
@@
However, since only approximately 3,200 PMTs were evaluated using the scanning station system, it is necessary to develop a method to estimate the PDE values at the seven zenith angles for the PMTs that were not tested in this system in order to constrain their non-uniform PC and ARC as well. A straightforward approach to predict these PDE values is to perform random sampling based on the measured data. To account for the potential dependence of PDE on zenith angles, a seven-dimensional PDE histogram is constructed using data collected with the scanning station system for each type of PMT, which includes 585 HPK PMTs, 1,719 NNVT normal-QE PMTs, and 939 NNVT high-QE PMTs. The random sampling of the PDE is carried out for each PMT using the seven-dimensional PDE histograms. A normalization factor is applied to the seven sampled values during the sampling process to ensure that the average sampled PDE of each PMT aligns with the PDE results obtained from the container system, where the average PDE is calculated by applying the weights presented in Table 1. This method allows the determination of PDE values at the seven reference positions through sampling for all PMTs that were not tested using the scanning station system.
@@
Since the PDE obtained above encompasses contributions from multiple positions on the PC in addition to the incident position, a detailed PMT optical simulation is conducted to correct for these additional contributions. Utilizing the PMT optical simulation software developed in [8], the detailed structure of the seven LED light sources has been newly implemented in the software, reflecting their real locations and orientations. This simulation allows for accurate accounting of the distributions of the AOI and the beam profile of light impinging on the PMT surface. Fig. 2 illustrates the zenith angle distribution contributing to the total absorption at 420 nm for HPK PMTs shown in red, NNVT normal-QE PMTs in green, and NNVT high-QE PMTs in blue. These results are obtained via the optical simulation for seven LED light sources, and the optical properties of the PC and ARC are derived from [8] with an assumption of uniform them. This figure indicates that the total absorption primarily arises from the incident position of the light beams, which is represented by the peak, and the ratio of the contributions from transmitted photons is depicted in the last panel and is used to correct the PDE by excluding these extra contributions based on their weights. It is observed that the two types of NNVT PMTs exhibit the similar distributions due to their same shape and structure, and compared to HPK PMTs, they gains significantly larger contributions from transmitted photons because of the considerably smaller electrodes within them, which reduces the loss of the transmitted light.
@@
Figure 2: The zenith angle distribution contributing to the total absorption for HPK PMTs shown in red, NNVT normal-QE PMTs in green, and NNVT high-QE PMTs in blue, and the ratio of the contributions from transmitted photons is shown in the last panel.
@@ 3.2 Reflectance data
The reflectance measurements were carried out before assembling the PMT covers, utilizing the setup developed in this work. In the first stage, a total of 565 20-inch PMTs were randomly selected for testing, which included 276 HPK PMTs, 115 NNVT normal-QE PMTs, and 174 NNVT high-QE PMTs. The reflectance was measured at 415 nm across four different zenith angles, including 0, 28, 51 and 72. In the second stage, only the PMTs tested with the scanning station system were selected, with measurements taken at four zenith angles and azimuth angle scans with a step of 45, involving 21 HPK PMTs, 38 NNVT normal-QE PMTs, and 45 NNVT high-QE PMTs.
The schematic diagram of the reflectance measurement setup is shown in Fig. 3. The system consists of an LED light source coupled with an optical fiber, with a central wavelength of 415 nm and a full width at half maximum of 14 nm. An integrating sphere, the ISP-50-8-REFL from Mikropack, is employed to collect the reflected light from the tested samples, along with a spectrometer and a computer for data storage, while a testing mold provides positional information on the PMTs. In the setup, optical fiber #1 transmits light from the LED to the integrating sphere, where the light beam is collimated using a lens integrated at the top of the sphere. The incident light beam exits the sphere at an AOI of 8 through a hole at the bottom, which has a diameter of approximately 8 mm. When a sample is placed over this hole, the reflected light is collected by the integrating sphere, and a fraction of this light is transmitted through optical fiber #2 to the spectrometer, where its intensity is measured as a function of wavelength. By comparing the measured intensity of the test sample with the reference sample of known reflectivity, the reflectance of the test sample can be determined. This measurement does not require dark conditions, as the background spectrum can be directly measured without samples placed over the hole of the integrating sphere. Fig. 4 illustrates an example of the measured spectra: the background (blue), the PMT (green), and the reference sample (orange).
@@
The systematic uncertainty has been carefully estimated, with the relative uncertainty determined to be approximately 2.5%. The primary contributions to this uncertainty arise from the LED stability and the reflectance uncertainty of the reference sample. Additionally, the reflectance of PMT was corrected using the ratio of the reflectance of the flat sample to the curved one for each zenith angle to eliminate the influence of the curvature of the PMT surface on the reflectance. Since no dependence of reflectance on the azimuthal direction was observed, the reflectance values measured at different azimuth angles are finally averaged for each zenith angle.
The measured results are presented in Fig. 5, showing data for all tested HPK PMTs (left), NNVT normal-QE PMTs (middle), and NNVT high-QE PMTs (right), and the solid blue lines represent the average reflectance for each type of PMT. Significant variations in reflectance can be observed across all types of PMTs. Specifically, the average reflectance at the zenith angle of zero degree for HPK PMTs is approximately 5.9%, which is lower than 10.8% and 16.4% observed for the NNVT normal-QE and NNVT high-QE PMTs, respectively. Additionally, the reflectance of HPK PMTs increases in areas closer to the equator; in contrast, the reflectance of NNVT PMTs exhibits a decreasing trend from 0 to about 50, followed by an increase.
@@
A similar random sampling approach to that used for estimating PDE is employed to assess the reflectance of PMTs that have not been measured. A four-dimensional reflectance histogram is constructed using the measured data for each type of PMT to account for the positional dependence of reflectance. Since the collected data indicate a weak correlation between reflectance and PDE, this relationship is not considered in the aforementioned sampling process.
@@ 4.1 Thickness map of the PC and ARC
$\chi^{2}=\sum_{i=1}^{4}\left(\frac{R(\theta_{i})-R_{theo}(d_{PC}(\theta_{i}),d_{ARC}(\theta_{i}))}{\sigma_{R}}\right)^{2}+$ (4.1)
$\sum_{j=1}^{5}\left(\frac{A(\theta_{j})-A_{theo}(d_{PC}(\theta_{j}),d_{ARC}(\theta_{j}))}{\sigma_{A}}\right)^{2}+$
$\left(\frac{CE-CE_{theo}}{\sigma_{CE}}\right)^{2}$
In this equation, $R(\theta_{i})$ represents the measured reflectance at the four zenith angles, while $R_{theo}$ and $A_{theo}$ are the calculated reflectance and absorptance, respectively, based on the optical theory. The predicted CE, $CE_{theo}$, is set to 0.93 for HPK PMTs and 1.0 for NNVT PMTs. The uncertainties $\sigma_{R}$, $\sigma_{A}$, and $\sigma_{CE}$ correspond to the measured reflectance, expected absorptance, and expected CE, respectively. Here, $\sigma_{CE}$ is assumed to be 3%, and $\sigma_{A}$ is estimated through error propagation from the measured PDE and its uncertainty, as the absorptance $A(\theta_{j})$ can be calculated using:
$A(\theta_{j})=\frac{PDE^{\prime}(\theta_{j})}{\rho\times CE}$ (4.2)
where $PDE^{\prime}$ represents the corrected PDE, excluding contributions from internal optical processes inside the PMTs. Only the PDE measured at five small zenith angles, namely from LED1 to LED5, are utilized in the fitting process. The value of the escape factor $\rho$ is at 420 nm from [8].
By minimizing Eq. 4.1, the free parameters $c_{0}$, $c_{1}$, and $c_{2}$ in Eq. 2.3 can be constrained for both the PC and ARC of each PMT, as well as the temporary CE. An example of the thickness curves derived from the fitting is illustrated in Fig. 6 for the HPK PMT (left), NNVT normal-QE PMT (middle), and NNVT high-QE PMT (right). The thickness of the ARC has been fixed to 49.1 nm for the NNVT normal-QE PMTs, due to its negligible effect on reflectance and absorptance, as demonstrated in Fig. 1. The obtained thickness distributions at the zenith angle of zero are presented in Fig. 7 for all HPK PMTs (left), NNVT normal-QE PMTs (middle), and NNVT high-QE PMTs (right).
@@
Figure 7: Thickness distributions of the PC and ARC at the zenith angle of zero degree for all HPK PMTs (left), NNVT normal-QE PMTs (middle), and NNVT high-QE PMTs (right).
@@
The aforementioned PMT optical simulation is conducted again for the seven LEDs in the scanning station system by incorporating the obtained thickness map for each PMT into the simulation. This allows us to determine the PDE contributions from various positions on the PC taking into account PC and ARC non-uniformity. From the simulation, both $a$ and $\beta$ can be derived for the seven reference positions. Using Eq. 2.2, the CE at these reference positions can be calculated for each PMT by solving the system of equations with the known PDE values at each point.
Fig. 8 illustrates the average CE across all PMTs of the same type. It is noted that some CE values are larger than one, which is a consequence of the assumption that the escape factor spectra remains unchanged for each type of PMT since decoupling CE from the escape factor based on the limited data exits challenges in this work.
@@ 4.3 Spectral response of optical properties
The refractive index and extinction coefficient was measured for the PC and ARC of each type of PMT in [8] within the wavelength range of 390 nm to 500 nm. In this study, the dispersion relation is utilized to extend the refractive index of the ARC, covering the wavelength range of interest for JUNO from 300 nm to 700 nm:
@@
where $p_{0}$ and $p_{1}$ are determined by fitting the measured $n$ values of the ARC between 390 nm and 500 nm. The results for the three types of PMTs are shown in Fig. 9(a).
For the PC, $n$ and $k$ values in the wavelength ranges of 300 nm to 390 nm and 500 nm to 700 nm are calculated using the permittivity results for the bialkali PC reported in [6] with the following formula:
@@
The calculated values in these two ranges are then normalized to the measured values at 390 nm and 500 nm as found in [8]. Fig. 9(b) illustrates the results for $n$ (solid lines) and $k$ (dashed lines) across the full wavelength range for the HPK PMT (red), NNVT normal-QE PMT (green), and NNVT high-QE PMT (blue).
The escape factor in the ranges of 300 nm to 390 nm and 500 nm to 700 nm is computed based on the measured quantum efficiency spectra in [3] and the respective absorption spectra, which is calculated using the $n$, $k$ and thickness values for both the PC and ARC in [8], and then the escape factors are normalized to align with the measured curve reported in [8]. The results for the escape factor are shown in Fig. 9(c) for the three types of PMTs.
Figure 9: (a) The refractive index of ARC as a function of wavelength for the HPK PMT (red), NNVT normal-QE PMT (green), and NNVT high-QE PMT (blue). (b) The $n$ (solid lines) and $k$ (dashed lines) of the PC as a function of wavelength for the HPK PMTs (red), NNVT normal-QE PMTs (green), and NNVT high-QE PMTs (blue). (c) The escape factor as a function of wavelength for the HPK PMT (red), NNVT normal-QE PMT (green), and NNVT high-QE PMT (blue).
4.4 Comparison of the average reflection and PDE
To evaluate the potential impacts of the comprehensive PMT optical model on the JUNO detector’s energy response, we compare the average reflectance and PDE derived from the comprehensive model to those obtained using the simplified PMT optical model, where all PMTs of the same type are assumed to share identical optical properties for the PC and ARC as described in [8].
Figure 10 illustrates the comparison results for the average reflectance (top) and PDE (bottom) at a zenith angle of zero degrees and a wavelength of 420 nm in water, as a function of the AOI, for the HPK PMT (left), NNVT normal-QE PMT (middle), and NNVT high-QE PMT (right). The blue solid lines represent the results with non-uniform PC and ARC, while the orange dashed lines correspond to the simplified model. The green solid lines in the lower panels show the ratio of the two models. Notably, the comprehensive model predicts an increase in reflections for the HPK PMTs, but a decrease of approximately 30% and 40% in the average reflectance for the NNVT normal-QE and NNVT high-QE PMTs, respectively. The comprehensive model has less than a 2% effect on the average PDE in regions of small AOI for all three PMT types, but a 5% to 10% impact on PDE at large AOI. It is important to note that the PDE values presented here do not include contributions from transmitted photons, which are event-dependent and can be accounted for through simulation.

### v2

@@ 3.1 PDE data
For each 20-inch PMT, the PDE is measured with a container system that employs a 420 nm light source to illuminate the entire PMT PC in air. Additionally, a subset of PMTs is evaluated for the PDE uniformity test with a scanning station system, which scans the PC with the seven LEDs that provide 420 nm light nearly perpendicular to the surface at the seven fixed zenith angles, meanwhile the LEDs can be rotated to cover the different azimuth angles. The PDE obtained from the container system is then corrected to align with the PDE as defined by the scanning station system, which averages the PDE values acquired at the positions of the seven LEDs, weighted by their respective surface areas, and Table 1 summarizes the surface area weights along with the zenith angles for each LED for both the NNVT and HPK PMTs. In this study, the PDE values averaged over the azimuthal angles for the seven zenith angles are utilized to extract the thickness maps of the PC and ARC.
@@
However, since only approximately 3,200 PMTs were evaluated with the scanning station system, it is necessary to develop a method to estimate the PDE values at the seven zenith angles for the PMTs that were not tested in this system in order to simulate their PDE non-uniformity as well. A straightforward approach to predict these PDE values is to perform random sampling based on the measured data. To account for the potential dependence of the PDE on zenith angles, a seven-dimensional PDE histogram is constructed according to data collected with the scanning station system for each type of PMT, which includes 585 HPK PMTs, 1,719 NNVT normal-QE PMTs, and 939 NNVT high-QE PMTs. The random sampling of the PDE is carried out with the seven-dimensional PDE histograms, which allows the determination of the PDE values at the seven reference positions for all the PMTs that were not tested with the scanning station system. A normalization factor is applied to the seven sampled values during the sampling process to ensure that the average sampled PDE of each PMT aligns with the PDE results acquired from the container system, where the average PDE is calculated by applying the weights presented in Table 1.
@@
Since the PDE obtained above encompasses the contributions from the transmitted photons absorbed at the multiple positions on the PC in addition to the direct absorption at the incident position, a detailed PMT optical simulation is conducted to distinguish the absorbed positions of the photons. Utilizing the PMT optical simulation software developed in [8], the detailed structure of the seven LED light sources has been newly implemented in the software, reflecting their real locations and orientations in the scanning station system, which allows for accurate accounting of the distributions of the AOI and the beam profile of light impinging on the PMT surface. The simulations for the seven LEDs were conducted with the uniform optical properties of the PC and ARC acquired from [8]. Fig. 2 illustrates the zenith angle distributions of the absorbed photons for the seven LEDs for the HPK PMTs shown in red, NNVT normal-QE PMTs in green, and NNVT high-QE PMTs in blue. This figure indicates that the total absorption primarily arises from the incident positions of the light beams, which are represented by the peak, while the rest shows the contributions from the transmitted photons. The ratio of the absorbed transmitted photons to the total absorbed photons is depicted in the last panel and applied to correct the PDE by excluding these extra contributions based on their weights. It is observed that the two types of the NNVT PMTs exhibit the similar distributions due to their same shape and structure, and compared to the HPK PMTs, they gain significantly larger contributions from the transmitted photons.
@@
Figure 2: The zenith angle distributions of the absorbed photons for the seven LEDs for the HPK PMTs shown in red, NNVT normal-QE PMTs in green, and NNVT high-QE PMTs in blue, and the ratio of the absorbed transmitted photons to the total absorbed photons is shown in the last panel.
@@ 3.2 Reflectance data
The reflectance measurements were carried out, utilizing the setup developed in this work. In the first stage, a total of 565 20-inch PMTs were randomly selected for testing, which included 276 HPK PMTs, 115 NNVT normal-QE PMTs, and 174 NNVT high-QE PMTs. The reflectance was measured at 415 nm across the four different zenith angles, including 0, 28, 51 and 72. In the second stage, only the PMTs tested with the scanning station system were selected, and measured at the four zenith angles as well as the azimuth angles scanned with a step of 45, involving 21 HPK PMTs, 38 NNVT normal-QE PMTs, and 45 NNVT high-QE PMTs.
The schematic diagram of the reflectance measurement setup is shown in Fig. 3. The system consists of an LED light source coupled with an optical fiber, with a central wavelength of 415 nm and a full width at half maximum of 14 nm. An integrating sphere, the ISP-50-8-REFL from Mikropack, is employed to collect the reflected light from the tested samples, along with a spectrometer and a computer for data storage, while a testing mold provides the positional information on the PMTs. In the setup, optical fiber #1 transmits light from the LED to the integrating sphere, where the light beam is collimated by a lens integrated at the top of the sphere. The incident light beam exits the sphere at an AOI of 8 through a hole at the bottom, which has a diameter of approximately 8 mm. When a sample is placed over this hole, the reflected light is collected by the integrating sphere, and a fraction of this light is transmitted through optical fiber #2 to the spectrometer, where its intensity is measured as a function of the wavelength. By comparing the measured intensity of the test sample with the reference sample of the known reflectance, the reflectance of the test sample can be determined. This measurement does not require dark conditions, as the background spectrum can be directly measured without samples placed over the hole of the integrating sphere. Fig. 4 illustrates an example of the measured spectra: the background (blue), the PMT (green), and the reference sample (orange).
@@
The systematic uncertainty has been carefully estimated, with the relative uncertainty determined to be approximately 2.5%. The primary contributions to this uncertainty arise from the LED stability and the reflectance uncertainty of the reference sample. Additionally, the reflectance of PMT was corrected by the ratio of the reflectance of the flat sample to the curved one for each zenith angle to eliminate the influence of the curvature of the PMT surface on the reflectance. Since no dependence of the reflectance on the azimuthal direction was observed, the reflectance values measured at the different azimuth angles are finally averaged for each zenith angle.
The measured results are presented in Fig. 5, showing data for all the tested HPK PMTs (left), NNVT normal-QE PMTs (middle), and NNVT high-QE PMTs (right), and the solid blue lines represent the average reflectance for each type of PMT. The significant variations in the reflectance can be observed across all the types of PMTs. Specifically, the average reflectance at the zenith angle of zero degree for the HPK PMTs is approximately 5.9%, which is lower than 10.8% and 16.4% observed for the NNVT normal-QE and NNVT high-QE PMTs, respectively. Additionally, the reflectance of the HPK PMTs increases in areas closer to the equator; in contrast, the reflectance of the NNVT PMTs exhibits a decreasing trend from 0 to about 50, followed by an increase.
@@
A similar random sampling approach to that applied for estimating the PDE is employed to assess the reflectance of PMTs that have not been measured. A four-dimensional reflectance histogram is constructed with the measured data for each type of PMT to account for the positional dependence of the reflectance. Since the collected data indicate a weak correlation between the reflectance and PDE, this relationship is not considered in the aforementioned sampling process.
@@ 4.1 Thickness map of the PC and ARC
$\chi^{2}=\sum_{i=1}^{4}\left(\frac{R(\theta_{i})-R_{theo}(d_{\mathrm{PC}}(\theta_{i}),d_{\mathrm{ARC}}(\theta_{i}))}{\sigma_{R}}\right)^{2}+$ (4.1)
$\sum_{j=1}^{5}\left(\frac{A(\theta_{j})-A_{theo}(d_{\mathrm{PC}}(\theta_{j}),d_{\mathrm{ARC}}(\theta_{j}))}{\sigma_{A}}\right)^{2}+$
$\left(\frac{\mathrm{CE^{\prime}}-\mathrm{CE}_{theo}}{\sigma_{\mathrm{CE^{\prime}}}}\right)^{2}$
In this equation, $R(\theta_{i})$ represents the measured reflectance at the four zenith angles, while $R_{theo}$ and $A_{theo}$ are the calculated reflectance and absorptance, respectively, based on the optical theory. The predicted $\mathrm{CE}_{theo}$ is set to 0.93 for the HPK PMTs and 1.0 for the NNVT PMTs. The uncertainties $\sigma_{R}$, $\sigma_{A}$, and $\sigma_{\mathrm{CE^{\prime}}}$ correspond to the measured reflectance, expected absorptance, and assumed constant CE’, respectively. Here, $\sigma_{\mathrm{CE^{\prime}}}$ is assumed to be 3%, and $\sigma_{A}$ is estimated through error propagation from the measured PDE and its uncertainty, as the absorptance $A(\theta_{j})$ can be calculated via:
$A(\theta_{j})=\frac{\mathrm{PDE}^{\prime}(\theta_{j})}{\rho\times\mathrm{CE^{\prime}}}$ (4.2)
where $\mathrm{PDE}^{\prime}$ represents the corrected PDE, which excludes the contributions from the absorbed transmitted photons to the total PDE. Only the PDE measured at the five small zenith angles, namely from LED1 to LED5, are utilized in the fitting process. The value of $\rho$ is at 420 nm from [8].
By minimizing Eq. 4.1, the free parameters $c_{0}$, $c_{1}$, and $c_{2}$ in Eq. 2.3 can be constrained for both the PC and ARC of each PMT. An example of the thickness curves derived from the fitting is illustrated in Fig. 6 for the HPK PMT (left), NNVT normal-QE PMT (middle), and NNVT high-QE PMT (right). The thickness of the ARC has been fixed to 49.1 nm for the NNVT normal-QE PMTs, due to its negligible effect on the reflectance and absorptance, as demonstrated in Fig. 1. The obtained thickness distributions at the zenith angle of zero are presented in Fig. 7 for all the HPK PMTs (left), NNVT normal-QE PMTs (middle), and NNVT high-QE PMTs (right).
@@
Figure 7: Thickness distributions of the PC and ARC at the zenith angle of zero degree for all the HPK PMTs (left), NNVT normal-QE PMTs (middle), and NNVT high-QE PMTs (right).
@@
The aforementioned PMT optical simulation is conducted again for the seven LEDs in the scanning station system by incorporating the constructed thickness map for each PMT into the simulation. This allows us to determine the contributions from the various positions on the PC to the PDE, taking into account the PC and ARC non-uniformity. From the simulation, both $a$ and $\beta$ can be derived for the seven reference positions. Applying Eq. 2.2, the CE at these reference positions can be calculated for each PMT by solving the system of equations with the PDE data at each point.
Fig. 8 presents the average CE across all the PMTs of the same type, illustrating the CE decreases with the increasing zenith angle. It is noted that some CE values are larger than one, which is a consequence of ignoring the variations of the escape factor among PMTs with the assumption that the escape factor spectra remains unchanged for each type of PMT, since decoupling the CE from the escape factor based on the limited data exits challenges in this work.
@@ 4.3 Spectral response of optical properties
The refractive index and extinction coefficient were measured for the PC and ARC of each type of PMT in [8] within the wavelength range of 390 nm to 500 nm. In this study, the dispersion relation is utilized to extend the refractive index of the ARC, covering the wavelength range of interest for JUNO from 300 nm to 700 nm:
@@
where $p_{0}$ and $p_{1}$ are determined by fitting the measured refractive index of the ARC between 390 nm and 500 nm. The results of the refractive index of the ARC for the three types of PMTs are shown in Fig. 9(a).
For the PC, the refractive index and extinction coefficient in the wavelength ranges of 300 nm to 390 nm and 500 nm to 700 nm are calculated with the permittivity ($\epsilon$) of the bialkali PC reported in [6] with the following formula:
@@
The calculated values in these two ranges are then normalized to the measured values at 390 nm and 500 nm as found in [8]. Fig. 9(b) illustrates the results for refractive index (solid lines) and extinction coefficient (dashed lines) across the full wavelength range for the HPK PMT (red), NNVT normal-QE PMT (green), and NNVT high-QE PMT (blue).
The escape factor in the ranges of 300 nm to 390 nm and 500 nm to 700 nm is computed based on the measured QE spectra in [3] and the respective absorption spectra, which is calculated with the aforementioned extended refractive index and extinction coefficient as well as the thickness values in [8] for both the PC and ARC, and then the escape factors are normalized to align with the measured curve reported in [8]. The results for the escape factor are shown in Fig. 9(c) for the three types of PMTs.
Figure 9: (a) The refractive index of the ARC as a function of the wavelength for the HPK PMT (red), NNVT normal-QE PMT (green), and NNVT high-QE PMT (blue). (b) The refractive index (solid lines) and extinction coefficient (dashed lines) of the PC as a function of the wavelength for the HPK PMTs (red), NNVT normal-QE PMTs (green), and NNVT high-QE PMTs (blue). (c) The escape factor as a function of the wavelength for the HPK PMT (red), NNVT normal-QE PMT (green), and NNVT high-QE PMT (blue).
4.4 Comparison of the average reflectance and PDE
To evaluate the potential impacts of the comprehensive PMT optical model on the JUNO detector’s energy response, we compare the average reflectance and PDE derived from the comprehensive model to those obtained utilizing the simplified PMT optical model, where all the PMTs of the same type are assumed to share identical optical properties for the PC and ARC as described in [8].
Fig. 10 illustrates the comparison results for the average reflectance (top) and PDE (bottom) at a zenith angle of zero degrees and a wavelength of 420 nm in water, as a function of the AOI, for the HPK PMT (left), NNVT normal-QE PMT (middle), and NNVT high-QE PMT (right). The blue solid lines represent the results with the non-uniform PC and ARC, while the orange dashed lines correspond to the simplified model. The green solid lines in the lower panels show the ratio of the two models. Notably, the comprehensive model predicts an increase in the reflectance for the HPK PMTs, but a decrease of approximately 30% and 40% in the average reflectance for the NNVT normal-QE and NNVT high-QE PMTs, respectively. The comprehensive model has less than a 2% effect on the average PDE in regions of the small AOI for all three PMT types, but a 5% to 10% impact on the PDE at the large AOI. It is important to note that the PDE values presented here do not include the contributions from the transmitted photons, which are event-dependent and can be accounted for through simulation.